OpenStudy (anonymous):

sin30 cos 60+sin 60 cos 30

OpenStudy (anonymous):

use Sin(a+b)=sin(90) =1

OpenStudy (anonymous):

sin(a+b)=sin a cos b + cos a sin b where a=30, b=60

OpenStudy (anonymous):

sinx cosy + siny cosx=sin(x+y) so sin(30+60)=sin90=1

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

;) yw